OpenStudy (anonymous):

problem is; lim y approaches 64 64-y/8-Square root of y not certain where to start/solve

OpenStudy (psymon):

Theres probably multiple ways, but I have a unique way of solving it that should work. We can actually do a difference of squares with the numerator. If you have anything minus anything, you can do this: \[(a-b) = (\sqrt{a} - \sqrt{b})(\sqrt{a}+\sqrt{b})\] So I will do exactly that with the top. We have 64 -y. So sqrt 64 is 8 and sqrt y is just rt(y). So I can break apart the top like: \[\frac{ (8-\sqrt{y})(8+\sqrt{y}) }{ 8-\sqrt{y} }\]Then you can go fromt here :3

OpenStudy (anonymous):

Thank youThank you!

OpenStudy (psymon):

Yep, np :3